OpenStudy (anonymous):

What is the radius of a circle with a circumference of 23.55 meters? Use 3.14 for pi and show your work.

OpenStudy (anonymous):

~4

OpenStudy (anonymous):

15/4

OpenStudy (lgbasallote):

divide 23.55 by 2pi

OpenStudy (lgbasallote):

\[C = 2\pi r\] \[r = \frac{C}{2\pi}\]

OpenStudy (lgbasallote):

if you have a calculator you'll get 3.75

OpenStudy (anonymous):

ok i see so look

OpenStudy (anonymous):

To find the radius of a circle with a circumference you would take 23.55 meters and use 3.14 for Pi. Divide = 23.55 / 3.14 = Radius =

OpenStudy (lgbasallote):

no no...you divide 23.55/2(3.14)

OpenStudy (anonymous):

okay

OpenStudy (lgbasallote):

you divide by 2 pi not just pi

OpenStudy (lgbasallote):

you get it right?

OpenStudy (anonymous):

ok yes